# Hessian-based lightweight neural network for brain vessel segmentation on a minimal training dataset

### v1

@@ 1 Introduction
There are very few such robotic systems for vascular surgery. R-One™ robotic assistance platform (Robocath) is a more recent innovation, specializing in robotic solutions for interventional cardiology, aiming to improve precision and safety in cardiovascular procedures [23].
@@ 1.1 Existing Datasets
Among the main datasets used for training and validating machine learning models for automatic segmentation, the IXI dataset stands out as particularly valuable. Unlike many benchmark datasets, IXI closely reflects the real-world conditions of clinical practice, where image quality and acquisition parameters vary significantly. It includes around 600 MRA scans of healthy adults, obtained using different MRI protocols: T1-, T2-, and PD-weighted imaging, magnetic resonance angiography (MRA), and diffusion-weighted imaging (15 directions). The data were collected at three London hospitals — Hammersmith Hospital (Philips 3T system), Guy’s Hospital (Philips 1.5T system), and the Institute of Psychiatry (GE 1.5T system).
@@ 1.2.2 Neural Networks for Segmentation
3D U-Net [11] established a foundation for volumetric medical imaging by extending the 2D U-Net architecture. Its encoder-decoder design, enhanced by skip connections, combines low- and high-level features for precise spatial localization. A weighted softmax loss enables training on sparsely annotated data—a common constraint in biomedical imaging. On-the-fly elastic deformations and batch normalization further improve robustness and convergence. While effective for general volumetric segmentation, its performance on thin tubular structures remains suboptimal due to inadequate topology preservation.
@@
The proposed model from paper [8] integrates advancements from prior work into a unified multi-task CNN. Three synergistic tasks—(1) voxel-wise segmentation, (2) vessel surface distance transform learning, and (3) centerline detection—are guided by a topology-aware loss that prioritizes structural fidelity over voxel-wise overlap. A decorrelation loss (DCL) regularizes the encoder to suppress scanner-specific features without adversarial training, enabling seamless multi-center generalization. Evaluated on six heterogeneous datasets, the model outperforms BRAVE-NET and VC-Net in DSC (0.74 vs. 0.71), Topological Coincidence (0.87 vs. 0.80), and pathology preservation ($80\%$ overlap for microaneurysms). A Python-based GUI supports real-time 3D visualization and interactive refinement, streamlining clinical workflows.

### v3

# Hessian-Based Lightweight Neural Network HessNet for State-of-the-Art Brain Vessel Segmentation on a Minimal Training Dataset

@@ 1 Introduction
There are very few such robotic systems for vascular surgery. R-One robotic assistance platform (Robocath) is a more recent innovation, specializing in robotic solutions for interventional cardiology, aiming to improve precision and safety in cardiovascular procedures [23].
@@ 1.1 Existing Datasets
Among the main datasets used for training and validating machine learning models for automatic segmentation, the IXI dataset stands out as particularly valuable. Unlike many benchmark datasets, IXI closely reflects the real-world conditions of clinical practice, where image quality and acquisition parameters vary significantly. It includes around 600 MRA scans of healthy adults, obtained using different MRI protocols: T1-, T2-, and PD-weighted imaging, magnetic resonance angiography (MRA), and diffusion-weighted imaging (15 directions). The data were collected at three London hospitals – Hammersmith Hospital (Philips 3T system), Guy’s Hospital (Philips 1.5T system), and the Institute of Psychiatry (GE 1.5T system).
@@ 1.2.2 Neural Networks for Segmentation
3D U-Net [11] established a foundation for volumetric medical imaging by extending the 2D U-Net architecture. Its encoder-decoder design, enhanced by skip connections, combines low- and high-level features for precise spatial localization. A weighted softmax loss enables training on sparsely annotated data a common constraint in biomedical imaging. On-the-fly elastic deformations and batch normalization further improve robustness and convergence. While effective for general volumetric segmentation, its performance on thin tubular structures remains suboptimal due to inadequate topology preservation.
@@
The proposed model from paper [8] integrates advancements from prior work into a unified multi-task CNN. Three synergistic tasks-(1) voxel-wise segmentation, (2) vessel surface distance transform learning, and (3) centerline detection are guided by a topology-aware loss that prioritizes structural fidelity over voxel-wise overlap. A decorrelation loss (DCL) regularizes the encoder to suppress scanner-specific features without adversarial training, enabling seamless multi-center generalization. Evaluated on six heterogeneous datasets, the model outperforms BRAVE-NET and VC-Net in DSC (0.74 vs. 0.71), Topological Coincidence (0.87 vs. 0.80), and pathology preservation ($80\%$ overlap for microaneurysms). A Python-based GUI supports real-time 3D visualization and interactive refinement, streamlining clinical workflows.